OpenStudy (anonymous):

Evaluate (5 + 2)2 − 40 + 3 = __________. I did it multiple times, and I either got 12, 3 or -6. help lol

Mehek (mehek14):

is it \((5+2)^2-40+3\)

OpenStudy (anonymous):

yes lol sorry about that

Mehek (mehek14):

\(5+2=7\\7^2=49\\49-40=9\\9+3=12\) just use PEMDAS

OpenStudy (anonymous):

Thank you lol. I figured ATLEAST one of mine were right.